OpenStudy (anonymous):

The area of a rectangle whose length is x + 4 and whose width is x – 9 .. is x2 – 5x – 36 ?

OpenStudy (anonymous):

Yes this is right..

OpenStudy (anonymous):

Area = Length x breadth .. => Area = (x+4)(x-9)

OpenStudy (anonymous):

\[Area = Length \times Width = (x+4)(x-9) = x^2 - 5x - 36\]

OpenStudy (precal):



OpenStudy (precal):

box multiplication